# ND-SDF: Learning Normal Deflection Fields for High-Fidelity Indoor Reconstruction

### v1

@@ A.3.3 Numerical Gradient
Following Neuralangelo Li et al. (2023), we employ numerical gradients w.r.t multi-resolution hash grids to enhance reconstruction quality. The numerical gradient is computed as follows (considering only the SDF value output of $f_{g}$):
@@ B.1 Ablation of partial unbiased rendering
In Section 3.5, we introduce an unbiased function. We partially apply this technique to thin structures indicated by the learned deflection angles, without doing so may result in an inability to converge and a decrease in reconstruction quality.
We attribute this issue to the inconsistency in perspective inherent in the transform function. According to TUVR Zhang et al. (2023):
$f^{\prime}_{g}(\textbf{r}(t_{i}))=\nabla f_{g}(\textbf{o}+t_{i}\textbf{v})\cdot v=\textbf{n}_{i}\cdot\textbf{v},$ (29)
where v represents the ray direction and $\textbf{n}_{i}$ denotes the normal at point $\textbf{x}_{i}$ (also $\textbf{r}(t_{i})$). We find that $\frac{\partial f_{g}(\textbf{r}(t_{i}))}{\partial t}$ is equivalent to the cosine of ray and normal.
Consider any other ray in space, denoted as $\textbf{r}^{\prime}(t)=\textbf{o}^{\prime}+t\textbf{v}^{\prime}$. Assuming that this ray also intersects at $\textbf{x}_{i}$ and samples the same point such that $\textbf{x}_{j}=\textbf{o}^{\prime}+t_{j}\textbf{v}^{\prime}=\textbf{x}_{i}$, the volumetric density of $\textbf{x}_{j}$ under the transform function is given by:
$\sigma(\textbf{r}^{\prime}({t_{j}}))=\alpha\Psi_{\beta}\left(\frac{-f_{g}(\textbf{r}^{\prime}({t_{j}}))}{\left|f_{g}^{\prime}(\textbf{r}^{\prime}({t_{j}}))\right|}\right),$ (30)
where $f_{g}(\textbf{r}(t_{i}))=f_{g}(\textbf{r}(t_{j}))$ because $\textbf{x}_{i}=\textbf{x}_{j}$ and $\textbf{n}_{i}=\textbf{n}_{j}$, but ${\left|f^{\prime}_{g}(\textbf{r}(t_{i}))\right|}\neq{\left|f^{\prime}_{g}(\textbf{r}^{\prime}(t_{j}))\right|}$ because $\textbf{n}_{i}\cdot\textbf{v}\neq\textbf{n}_{j}\cdot\textbf{v}^{\prime}$. The inconsistency in perspective leads to ambiguity in the volume density at the same sampling point, causing complex indoor surfaces to fail to converge. Therefore, to reduce this ambiguity, we choose to apply the transform function only to fine regions indicated by the deflection angle.
The quantitative ablation results for partial unbiased rendering are presented in Table 8, where the improved F-score underscores the significance of this strategy.
@@
In Figure 8, qualitative comparisons on ScanNet are presented. As indicated by the blue box, the reconstructed surfaces exhibit wrinkles and irregularities without the application of the proposed partial unbiased rendering strategy and fail to capture the fine structural details of the chair legs highlighted in the red box. After applying the strategy, the overall surface becomes precise and smooth. These visual results demonstrate the superiority of the proposed partial unbiased rendering method.
Method F-score
w/o partial 0.662
w partial 0.686
Table 8: Ablation of Partial Unbiased Rendering. We observed a decrease in reconstruction quality after omitting partial unbiased rendering.
@@ B.2 Ablation of adaptive monocular depth loss
Notably, we use the same method to adjust depth supervision (Eq. 13) based on deflection angle. This simplified treatment rests on a strong foundation for two reasons. First, both the depth and normal cues originate from the same pretrained model Eftekhar et al. (2021) and share a common inductive bias. Second, monocular depth lacks scale and cannot provide accurate supervision. In implicit surface reconstruction with priors, depth cues are primarily utilized to facilitate convergence and accelerate surface formation. Automatically and continuously masking high-frequency areas based on the deflection angle has no adverse impact on the recovery of details.
@@
Method F-score
w/o $\mathcal{L}^{d}_{depth}$ 0.664
w $\mathcal{L}^{d}_{depth}$ 0.686
Table 9: Ablation of adaptive depth prior loss. We observed a decrease in reconstruction accuracy after omitting the adaptive depth loss.
@@ C.2 ScanNet++
We present detailed quantitative results of the selected six scenes from ScanNet++ in Table 10 and provide more visualization results in Figure 10. Compared to the baselines, our method yields significantly improved reconstructions both visually and quantitatively. Notably, our method contributes substantially to capturing thin and fine-grained structures.
@@ C.3 Tanks and temples
We present additional qualitative comparisons with previous state-of-the-art baselines, including MonoSDF Yu et al. (2022b) and NeuRIS Yu et al. (2022b), as depicted in Figure 11. Notably, ND-SDF achieves accurate dense surfaces replete with extensive details.
As illustrated in Figure 12 and Figure 13, we further extracted the reconstructed mesh at a resolution of 2048. The mesh, with this resolution, demonstrates that our method significantly enhances the generation of small and complex structures, even in challenging large-scale indoor scenes. The hash grid for all experiments remains at the default settings, with resolutions spanning from $2^{5}$ to $2^{11}$ across 16 levels, and the size of hashmap staying at $2^{19}$.
@@ C.4 More information about dataset
The qualitative results on ScanNet in the main content demonstrate that our method can capture fine details, such as the chair legs, which are even missed by the ground truth (GT). This highlights a known issue regarding the reliability of ScanNet GT meshes. The ScanNet dataset, constructed in 2017, is based on RGB-D data captured by an iPad with an additional structure sensor. The images have relatively low resolution and suffer from motion blur and lighting variations. In addition, due to the limitations of the bundle-fusion algorithm, the GT surfaces lack fine structures, which can lead to a decrease in evaluation performance in these areas. This also explains why we tend to conduct quantitative ablation studies on ScanNet++, which provides more reliable evaluation results.
Figure 9: Visualization results on Replica dataset.
Figure 10: Qualitative Comparison on ScanNet++. We compare ND-SDF with several baseline methods, including VolSDF, Baked-Angelo, and MonoSDF. Notably, only MonoSDF leveraged monocular priors. Our method facilitates both smoothness and details, while MonoSDF loses substantial fine structures, and Baked-Angelo cannot handle textureless smooth regions such as floors.

### v2

@@ A.3.3 Numerical Gradient
Following Neuralangelo Li et al. (2023), we employ numerical gradients w.r.t multi-resolution hash grids to promote surface consistency, as proved effective in their work. The numerical gradient is computed as follows (considering only the SDF value output of geometry network $f_{g}$):
@@
This term decreases exponentially during training, similar to Neuralangelo. In fact, our tests indicate that it has minimal impact on surface smoothness, resulting in only a slight improvement compared to its absence. We attribute this to the use of monocular cues, as the cues are particularly accurate in smooth regions.
@@ B.1 Ablation of partial unbiased rendering
In Section 3.5, we introduced an unbiased function. We partially apply this technique to thin structures indicated by the learned deflection angles, without doing so may result in an inability to converge and a decrease in reconstruction quality.
We attribute this issue to the inconsistency in perspective inherent in the transform function. We compute the derivative of $f_{g}$ (considering only its SDF value output) w.r.t the sampling distance $t_{i}$ as follows:
$f^{\prime}_{g}(\textbf{r}(t_{i}))=\nabla f_{g}(\textbf{o}+t_{i}\textbf{v})\cdot\textbf{v}=\textbf{n}_{i}\cdot\textbf{v},$ (30)
where v is the ray direction and $\textbf{n}_{i}$ represents the geometry normal at point $\textbf{x}_{i}$. It means that $\frac{\partial f_{g}(\textbf{r}(t_{i}))}{\partial t_{i}}$ is equivalent to the cosine of the ray direction and the normal.
Consider another ray in space, denoted as $\textbf{r}^{\prime}(t)=\textbf{o}^{\prime}+t\textbf{v}^{\prime}$. Assuming that this ray also passes through $\textbf{x}_{i}$, sampling the same point such that $\textbf{x}_{j}=\textbf{o}^{\prime}+t_{j}\textbf{v}^{\prime}=\textbf{x}_{i}$. According to Eq. 15, the volumetric densities of $\textbf{x}_{i}$ and $\textbf{x}_{j}$ under the transform function are given by:
$\sigma(\textbf{r}({t_{i}}))=\frac{1}{\beta}\Psi_{\beta}\left(\frac{-f_{g}(\textbf{r}({t_{i}}))}{\left|f_{g}^{\prime}(\textbf{r}({t_{i}}))\right|}\right),\sigma(\textbf{r}^{\prime}({t_{j}}))=\frac{1}{\beta}\Psi_{\beta}\left(\frac{-f_{g}(\textbf{r}^{\prime}({t_{j}}))}{\left|f_{g}^{\prime}(\textbf{r}^{\prime}({t_{j}}))\right|}\right).$ (31)
We observe that $f_{g}(\textbf{r}(t_{i}))=f_{g}(\textbf{r}^{\prime}(t_{j}))$ because $\textbf{x}_{i}=\textbf{x}_{j}$, but ${\left|f^{\prime}_{g}(\textbf{r}(t_{i}))\right|}\neq{\left|f^{\prime}_{g}(\textbf{r}^{\prime}(t_{j}))\right|}$ because $\textbf{n}_{i}\cdot\textbf{v}\neq\textbf{n}_{j}\cdot\textbf{v}^{\prime}$ even though $\textbf{n}_{i}=\textbf{n}_{j}$. The inconsistency in perspective leads to ambiguity in the volume density at the same sampling point, causing complex indoor surfaces to fail to converge. Therefore, to reduce this ambiguity as much as possible, we choose to apply the transform function only to fine regions indicated by the deflection angle.
The quantitative ablation results on ScanNet for partial unbiased rendering are presented in Table 8, where the improved F-score highlights the effectiveness of this strategy.
Figure 10 provides qualitative comparisons. As indicated by the blue box, the surfaces reconstructed without the proposed partial unbiased rendering exhibit wrinkles and irregularities, and the fine structural details of the chair legs (highlighted in red) are not well captured. With the application of this strategy, the surfaces become significantly smoother and more precise. These visual results underscore the advantages of the partial unbiased rendering approach.
@@
Method w/o partial w partial
F-score 0.707 0.746
Table 8: Ablation of Partial Unbiased Rendering on ScanNet. We observed a decrease in reconstruction quality after omitting partial unbiased rendering.
@@ B.2 Ablation of adaptive monocular depth loss
Notably, we apply the same method to adjust depth supervision (Eq. 13) based on the deflection angle. This simplified treatment rests on a strong foundation for two reasons. First, both depth and normal cues share a common inductive bias, as they stem from the same pretrained model Eftekhar et al. (2021). Second, due to the inherent lack of scale in monocular depth, it cannot provide precise supervision. In implicit surface reconstruction with priors, depth cues are primarily utilized to facilitate convergence and accelerate surface formation. Therefore, automatically masking high-frequency areas based on the deflection angle has no adverse impact on the recovery of details.
@@
Method w/o $\mathcal{L}^{d}_{depth}$ w $\mathcal{L}^{d}_{depth}$
F-score 0.664 0.686
Table 9: Ablation of adaptive depth prior loss on ScanNet++. We observed a decrease in reconstruction accuracy after omitting the adaptive depth loss.
Figure 11: Visualization results on Replica dataset.
@@ C.2 ScanNet++
We present detailed quantitative results of the selected six scenes from ScanNet++ in Table 10 and provide more visualization results in Figure 12. Compared to the baselines, our method yields significantly improved reconstructions both visually and quantitatively, notably excelling at capturing thin, fine-grained structures while maintaining smoothness in flat regions.
@@ C.3 Tanks and temples
We present additional qualitative comparisons with previous state-of-the-art baselines, including MonoSDF Yu et al. (2022b) and NeuRIS Yu et al. (2022b), as depicted in Figure 13. It is evident that ND-SDF achieves accurate dense surfaces replete with extensive details.
As illustrated in Figure 14 and Figure 15, we further extracted the reconstructed meshes at a resolution of 2048. At this high resolution, the meshes demonstrate that our method significantly enhances the generation of small and thin structures, even in challenging large-scale indoor scenes. For all experiments, the hash grid settings were kept at their defaults, with resolutions ranging from $2^{5}$ to $2^{11}$ across 16 levels, and a hashmap size of $2^{19}$.
@@ C.4 More information about dataset
The qualitative results on ScanNet presented in the main content (as shown in Figure 4) demonstrate that our method captures fine details, such as the chair legs, which are even overlooked by the ground truth (GT). This highlights a known issue regarding the reliability of ScanNet GT meshes. Constructed in 2017, the ScanNet dataset is based on RGB-D data captured by an iPad with an additional structure sensor. The images suffer from relatively low resolution, motion blur, and lighting variations. Furthermore, the limitations of the bundle-fusion algorithm result in GT surfaces lacking fine structures, which can adversely affect evaluation performance in these areas. This underscores our preference for conducting quantitative ablation studies on ScanNet++, which provides more reliable evaluation results.
Figure 12: Qualitative Comparison on ScanNet++. We compare ND-SDF with several baseline methods, including VolSDF, Baked-Angelo, and MonoSDF. Notably, only MonoSDF and Ours leverage monocular priors. Our method effectively balances smoothness and detail, whereas MonoSDF loses significant fine structures, and Baked-Angelo struggles to handle textureless smooth regions, such as floors.